OpenStudy (anonymous):

Can somebody please help me? I don't understand the lesson and how to work these problems out. Thank you! Given the function ƒ(x ) = 3x - 14, find ƒ(3). -5 9 5

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

\[\color{blue}{f(x) = 3x - 14}\]\[f(?) = 3? - 14\]\[f(\heartsuit ) = 3\heartsuit - 14\]\[f(1) = 3(1) - 14 = 3 - 14 = -11\]Remember: you just replace all \(x\)'s.

OpenStudy (anonymous):

But what do you replace the x's with?

OpenStudy (ccswims):

replace the "x" with 3, because f(3) is saying that now we know hat x is.

OpenStudy (ccswims):

what*

OpenStudy (anonymous):

Oh! Yeah, now that makes sense :) Thank you, @ccswims and @ParthKohli :)

Parth (parthkohli):

No problem.

OpenStudy (anonymous):

:) I wish I could give out 2 medals :) You were both helpful, I think I can do this lesson on my own now!